# Energy-based Contact Planning under Uncertainty for Robot Air Hockey

### v1

@@ Abstract
Planning robot contact often requires reasoning over a horizon to anticipate outcomes, making such planning problems computationally expensive. In this letter, we propose a learning framework for efficient contact planning in real-time subject to uncertain contact dynamics. We implement our approach for the example task of robot air hockey. Based on a learned stochastic model of puck dynamics, we formulate contact planning for shooting actions as a stochastic optimal control problem with a chance constraint on hitting the goal. To achieve online re-planning capabilities, we propose to train an energy-based model to generate optimal shooting plans in real time. The performance of the trained policy is validated in simulation and on a real-robot setup. Furthermore, our approach was tested in a competitive setting as part of the NeurIPS 2023 Robot Air Hockey Challenge.
@@ I Introduction
Planning and control through non-prehensile contacts is an essential skill for robots to interact with their environment. Model-based approaches enable robots to anticipate the outcome of contact interactions given a candidate action allowing them to find an action with the desired outcome. While model-based planning approaches are shown to be successful at generating contact-rich plans for slow tasks [1, 2], highly dynamic tasks require the agent to regenerate contact plans at a sufficiently high rate for reacting to inherent perturbations. Highly dynamic tasks not requiring reasoning through contacts have historically been used as a testbed for hardware and algorithms in robotics. These tasks include different types of games and sports, such as ball-in-a-cup [3, 4], juggling [5, 6], diabolo [7]. Dynamic tasks involving contacts, such as soccer [8], tennis [9], table tennis [10, 11], and air hockey [12, 13], are typically approached with reinforcement learning methods to off-load the computationally expensive reasoning through contacts to an offline exploration phase. Yet, these tasks have in common that contacts with the ball or puck are instantaneous, i.e. the contact happens in a short period of time, resulting in a jump in the state of the object. The reasoning over the contact between the robot and the object of interest can therefore be divided into three segments of the planning horizon: i) Moving the robot into contact, ii) the contact itself at a single time instance, and iii) the passive trajectory of the object after contact.
Figure 1: The proposed control framework enables a robot arm to autonomously play matches of air hockey. First, a learned stochastic model of contact dynamics is used to predict the trajectory of a puck. An energy-based contact planner is then trained to generate agile behavior in real time.
In this letter, we exploit the separability in the planning horizon by combining a model-based control approach for moving the robot into contact with a learning-based approach for planning the next best contact that results in the desired trajectory of the object. Towards this end, we learn a mixture of linear-Gaussian modes for modeling the object dynamics from data, which allows us to extract a stochastic model for the contact between the robot and the object. Based on the learned model, we train an energy-based contact policy by generating example contacts that are optimal w.r.t. a stochastic optimal control objective offline. During the online phase, we retrieve optimal contact plans from the energy-based policy using derivative-free inference in real time.
@@
In the following, we present our approach in the context of the highly-dynamic game air hockey. Fig. 2 illustrates the online control framework that consists of state estimation, the proposed learning-based contact planner (shooting policy), and a subsequent model-based robot controller (MPC). For the robot controller, we use a sampling-based model-predictive controller [14] that enforces the execution of the contact plan while respecting safety constraints such as collision avoidance with the walls. We summarize our key scientific contributions as follows:
@@
  • We present an approach for learning the parameters of a stochastic model for discontinuous contact dynamics as a mixture of linear-Gaussian modes.
@@
  • We formulate the planning of contacts as a chance-constrained stochastic optimal control problem.
  • We propose an approach for training an energy-based model to capture the optimal policy according to the chance-constrained stochastic optimal control problem.
Our approach is experimentally validated in a dynamic air hockey shooting task with comparisons to control-based and reinforcement learning baselines. We additionally deploy our framework in a competitive setting as part of the NeurIPS 2023 Robot Air Hockey Challenge (see Fig. 1). Our framework outperforms all other approaches in real-robot matches, establishing a new state-of-the-art in robot air hockey.
@@
Figure 2: Overview of the interplay between puck state estimation $\bullet$ and robot control $\bullet$ in our framework for agile robot air hockey. The contact planner uses the estimated puck state to predict the puck trajectory based on the learned model. It subsequently plans a shooting angle that is used to construct an optimal control objective solved within a model-predictive controller. All modules are updated at a control rate of $50$ Hz.
@@ II Related Work
Another use of the robot air hockey setting is as a testbed for learning algorithms. In [26], deep reinforcement learning techniques are used to learn on planar robots, while in [12], both the planar 3-DoF and the 7-DoF cobot air hockey tasks are used to learn control policies in simulation. More recent techniques directly use the real 7-DoF air hockey setting as a testbed for learning algorithms: in [27], the authors use learning-to-plan techniques to generate air hockey hitting trajectories in the real-world setting, while  [13], this task is used to perform real-world reinforcement learning.
In general, existing solutions to the robot air hockey problem can be categorized in two main directions: learning-based approaches [17, 26, 13], and control-based approaches [21, 23, 24]. Generally speaking, pure control-based approaches lead to better and faster solutions than learning-based methods but require considerable efforts in engineering and model identification, and are particularly challenging to implement to run at real-time control rates. Instead, pure learning-based approaches obtain a worse-quality solution but make it possible to obtain more robust behaviors by relying on domain randomization and fine-tuning on the real platform. In this letter, we aim to combine the advantages of learning-based and control-based approaches. We exploit the optimality of control-based approaches for controlling the robot without consideration of the puck and exploit the robustness and flexibility of learning-based approaches to efficiently generate plans for the contact between the robot and the puck to maximize the chance of scoring.
III Learning Stochastic Contact Models
Planning and controlling the contacts of the robot with the puck requires the anticipation of puck trajectories. To enable this, we learn a simplified stochastic model of the puck dynamics for i) estimating the current state of the puck online, ii) predicting the trajectory of the puck online, and iii) solving a stochastic optimal control problem to plan the next best contact between the robot and the puck.
@@ III-A Mixture of linear-Gaussian Contact Dynamics
Suppose that $\bm{x}_{k}^{p}\in\mathbb{R}^{2}$ is the position of the puck w.r.t. the surface of the air hockey table at time step $k$. The robot interacts with the puck by making contact with its mallet, i.e. the circular part of the robot’s end-effector. The position of the mallet is denoted with $\bm{x}_{k}^{m}\in\mathbb{R}^{2}$ w.r.t. the surface of the air hockey table. We assume that the robot arm is controlled such that the mallet maintains contact with the table at all times. In order to efficiently perform rollouts of the puck dynamics, we impose a piecewise-linear structure on the model. Fig. 3 illustrates the three modes that we present in the following: 1) Floating, 2) Puck-Wall Collision, and 3) Puck-Mallet Collision. To account for modeling errors introduced through the piecewise-linear structure, we model each mode as a conditional Gaussian distribution, resulting in a mixture of linear-Gaussian contact dynamics. In the following, we present the individual modes and their respective parameters that are learned subsequently. Note that the learnable parameters define the stochastic prediction of the puck velocity, while the one-step prediction of the position is derived from numerical integration and is deterministic.
Figure 3: Illustration of three modes of the puck dynamics that are parameterized as linear-Gaussian models. Mode 1) captures the dynamics of the puck $\bullet$ when floating on the surface of the table. Mode 2) captures collisions between puck and walls. Mode 3) models collisions between puck and the mallet $\bullet$ in a contact-aligned frame $\mathcal{C}$. The parameters for the nominal dynamics and the corresponding uncertainty are learned from data.
@@ III-B Learning Model Parameters from Data
Given the parameters of the joint probability distribution, the parameters of the linear-Gaussian models can be computed by conditioning the probability distribution on the input $\bm{\xi}$. The parameters are thus given by
@@ III-D Probability of Hitting the Goal
Fig. 4 illustrates stochastic rollouts for various initial conditions of a shot. Evaluating $\hat{\mathrm{G}}$ as defined in (11), we observe that those initial conditions have a significant effect even if the expected puck trajectory hits the center of the goal for all conditions. Fast shots (Fig. 4-b) accumulate less uncertainty compared to slow shots (Fig. 4-a) due to the fact that the modeled process noise is constant over time. Compared to bank shots, direct shots accumulate less uncertainty during rollout since collisions with a wall add significant process noise (Fig. 4-c).
IV Fast Contact Planning under Uncertainty
@@
The learned dynamics model enables the prediction of uncertain puck trajectories for contact planning. In particular, we aim to find contact states of the mallet that result in desired puck trajectories after contact. The proposed contact planning module is based on stochastic optimal control, optimizing for the mallet state at contact. We combine the optimization of the contact state with a model-based robot controller that drives the robot to the desired contact state at the desired contact time (cf. Fig. 2).
@@ IV-B Shooting Angle as Reduced Action Space
IV-C Training an Energy-based Shooting Policy
The long-horizon predictions required to plan shooting actions make it difficult to operate at a rate that is sufficient for agile behavior. We address this challenge by training an energy-based model to reproduce the solutions to the stochastic optimal shooting problem (12) in real-time. Due to implicit policy representation, energy-based models are particularly well-suited for multi-modal solution spaces [28]. In the case of shooting, different modes of the policy include straight shots, single bank shots, and double bank shots.
We train the energy-based model by solving the computationally expensive shooting angle optimization offline and using the results as training data. Namely, we first generate a dataset of shooting angles for $N$ different scenarios, i.e. puck states at the time of contact $\{\bm{s}_{i}\}_{i=1}^{N}$. Due to the one-dimensional parametrization of the action space, we efficiently explore the space of shooting angles for each initial puck state by sampling $M$ candidate shooting angles $\{u_{i}^{j}\}_{j=1}^{M}$. We subsequently compute the stochastic rollout of the puck trajectory as presented in Sec. III-D and evaluate the objective and chance constraint from (12). The best-performing sample $\hat{u}_{i}$ is then used as a positive example for training, and the remaining $M-1$ samples as negative counter-examples. We finally obtain a dataset of $M\times N$ state-action pairs $\{\bm{s}_{i},\hat{u}_{i},\{u_{i}^{j}\}_{j=1}^{M-1}\}_{i=1}^{N}$ and train the energy model $E_{\theta}(\bm{s},u)$ using an InfoNCE-style [29] loss
$\mathcal{L}_{\text{InfoNCE}}$ $=\sum_{i=1}^{N}-\log\left(\tilde{p}_{\theta}\big{(}\hat{u}_{i}|\bm{s}_{i},\{u^{j}_{i}\}_{j=1}^{M-1}\big{)}\right),$ (14)
where $\tilde{p}_{\theta}\big{(}\hat{u}_{i}|\bm{s}_{i},\{u^{j}_{i}\}_{j=1}^{M-1}\big{)}$ represents a likelihood with
@@ IV-D Online Inference with Warm-Starting
To solve (12) given the estimated puck state $\hat{\bm{s}}_{0^{-}}$, we search for the state-action pair that minimizes energy, i.e.
@@
For real-time optimization, we leverage direct access to the learned energy landscape of the EBM by executing sampling iterations concurrently with other components of the control loop. This allows us to simultaneously refine contact plans as trajectories are being executed on the robot. The online retrieval of optimal shooting angles is based on the derivative-free optimization procedures used in [28]. As in the offline scenario, we initiate an online shooting action by uniformly sampling $N$ candidate actions. Based on the corresponding energies, candidates are resampled with replacement to warm-start optimization at each following timestep. To converge towards a solution with minimum implicit energy, reductions to the sampling scale $\sigma$ are applied at each timestep, keeping the optimal contact angle $\hat{u}$ as a reference for the mid-level trajectory planner. A full iteration of the EBM optimization is outlined in Alg. 1. We observe that the learned energy models and utilized optimization procedure efficiently retrieve multimodal contact plans to produce desired behaviors, as shown in Fig. 5.
@@ V-A2 EBM Architecture and Training
The energy-based shooting model consists of a multilayer perceptron with $2$ hidden, fully connected layers of $128$ neurons each. The model is trained on $N\!=\!3000$ initial puck states, with $M\!=\!100$ action samples. Training required between $500$ and $1000$ epochs to converge for satisfactory performance using the Adam [30] optimizer with a decaying learning rate.
@@ V-B Experimental Setup
The experiment is conducted with a KUKA iiwa14 LBR manipulator equipped with a mallet end-effector that is attached to a passive joint for seamless contact with the table surface. Experiments are carried out in a simulated MuJoCo environment and on a real-world setup (cf. Fig. 1) for evaluation of sim2real transfer. We evaluate three instances of the proposed approach by using different parameters for the chance-constrained optimization problem in (12). Ours #1: a conservative policy that prioritizes accuracy ($\lambda_{1}\!=\!1,\lambda_{2}\!=\!0,\beta\!=\!0.5$); Ours #2: a balanced policy that compromises between accuracy and puck speed ($\lambda_{1}\!=\!1,\lambda_{2}\!=\!0.2,\beta\!=\!0.5$); and Ours #3: an aggressive policy that prioritizes puck speed ($\lambda_{1}\!=\!0,\lambda_{2}\!=\!1,\beta\!=\!0.5$). Simulated results are also illustrated in Fig. 5 for initial puck velocities of zero. We compare the three instances of our contact planner with: CB, a baseline that utilizes conventional planning and control methods [24]; and ATACOM, a reinforcement learning approach for learning a robot policy [12]. Note that the ATACOM policy is trained in simulation and deployed in the physical experiment without additional tuning or retraining.

### v2

# Distilling Contact Planning for Fast Trajectory Optimization in Robot Air Hockey

@@ Abstract
Robot control through contact is challenging as it requires reasoning over long horizons and discontinuous system dynamics. Highly dynamic tasks such as Air Hockey additionally require agile behavior, making the corresponding optimal control problems intractable for planning in realtime. Learning-based approaches address this issue by shifting computationally expensive reasoning through contacts to an offline learning phase. However, learning low-level motor policies subject to kinematic and dynamic constraints can be challenging if operating in proximity to such constraints is desired. This paper explores the combination of distilling a stochastic optimal control policy for high-level contact planning and online model-predictive control for low-level constrained motion planning. Our system learns to balance shooting accuracy and resulting puck speed by leveraging bank shots and the robot’s kinematic structure. We show that the proposed framework outperforms purely control-based and purely learning-based techniques in both simulated and real-world games of Robot Air Hockey.
@@ I Introduction
Planning and control through non-prehensile contacts is an essential skill for robots to interact with their environment. Model-based approaches enable robots to anticipate the outcome of contact interactions given a candidate action, allowing them to find an action with the desired outcome. Although model-based planning approaches have been shown to be successful in generating contact-rich plans for slow tasks [24, 13], highly dynamic tasks require the agent to regenerate contact plans at a sufficiently high rate to react to inherent perturbations. These tasks have historically been used as a testbed for hardware and algorithms in robotics, with different types of games and sports, such as ball-in-a-cup [14, 17], juggling [26, 25], and diabolo [31]. Dynamic tasks that involve contact, such as soccer [9], tennis [32], table tennis [22, 5], and air hockey [19, 20], are typically approached with reinforcement learning methods to off-load computationally expensive reasoning through contacts to an offline exploration phase. Yet, these tasks have in common that contact with the ball or puck is instantaneous, resulting in a jump in the object state. Reasoning about the contact between the robot and the object of interest can therefore be divided into three segments of the planning horizon: i) Moving the robot into contact, ii) the contact itself at a single time instance, and iii) the passive trajectory of the object after contact.
Figure 1: The proposed control framework enables our robot to autonomously play matches of air hockey. The dynamic game requires the robot to predict puck trajectories, plan the best contact, and coordinate its joints to generate high velocities without hitting a wall or lifting the mallet from the table.
This paper exploits such separability in the highly dynamic game of air hockey (Fig. 1) by combining a learning-based approach for contact planning with model-based control for moving the robot into contact. We show that distilling a stochastic optimal control policy enables us to effectively reduce the planning horizon required to generate desired behaviors, allowing our agent to operate in real-time. Furthermore, we highlight the interpretability of our approach by producing various shooting behaviors through different formulations of the optimal control cost, while exploiting the kinematic structure of the robot.
@@
Fig. 2 illustrates the online control framework that consists of state estimation, a learning-based contact planner (shooting policy), and a subsequent model-based robot controller (MPC). For state estimation and prediction, we learn a stochastic model of contact between the robot and the puck from data as a mixture of linear-Gaussian modes. Based on the learned model, we generate a dataset of example contacts that are optimal w.r.t. a stochastic optimal control objective, and distill the resulting policy through behavioral cloning. During the online phase, we retrieve optimal contact plans from the distilled policy using derivative-free inference in realtime. Finally, for the robot controller, we use a sampling-based model-predictive controller [12] that enforces the execution of the contact plan while respecting safety constraints such as collision avoidance with the walls. We summarize our key contributions as follows:
@@
  • We present an approach for learning the parameters of a stochastic model for discontinuous contact dynamics in robot air hockey as a mixture of linear-Gaussian modes.
@@
  • We formulate contact planning for robot air hockey as a chance-constrained stochastic optimal control problem.
  • We propose an approach for distilling an optimal contact policy by training an implicit model to allow for planning in real-time.
After presenting our technical contributions, we provide experimental comparisons to control-based and reinforcement learning baselines. Our approach has furthermore outperformed all other approaches tested in a competitive setting [21].
@@
Figure 2: Overview of the interplay between puck state estimation $\bullet$ and robot control $\bullet$ for closed-loop agile robot air hockey. The contact planner uses the estimated puck state to predict the puck trajectory based on the learned model. It subsequently plans a shooting angle that is used to construct an optimal control objective solved within a model-predictive controller. Robot trajectories are computed at a control rate of 50 Hz.
@@ II Related Work
Another use of the robot air hockey setting is as a testbed for learning algorithms. In [29], deep reinforcement learning techniques are used to learn on planar robots, while in [19], both the planar 3-DoF and the 7-DoF cobot air hockey tasks are used to learn control policies in simulation. More recent techniques directly use the real 7-DoF air hockey setting as a testbed for learning algorithms: in [15], the authors use learning-to-plan techniques to generate air hockey hitting trajectories in a real-world setting, while in  [20], this task is used to perform real-world reinforcement learning.
In general, existing solutions to the robot air hockey problem can be categorized in two main directions: learning-based approaches [3, 29, 20], and control-based approaches [28, 1, 18]. Generally speaking, pure control-based approaches lead to better and faster solutions than learning-based methods but require considerable efforts in engineering and model identification, and are particularly challenging to implement and run at realtime control rates. Instead, pure learning-based approaches obtain a lower-quality solution but make it possible to obtain more robust behaviors by relying on domain randomization and fine-tuning on the real platform. We aim to combine the advantages of learning-based and control-based approaches. We exploit both the optimality of control-based approaches for controlling the robot without considering the puck and the robustness and flexibility of learning-based approaches to efficiently generate plans for the contact between the robot and the puck to maximize the scoring probability.
III Learning a Stochastic Contact Model
Planning and controlling the contacts of the robot with the puck requires the anticipation of puck trajectories before and after contact. To enable this, we learn a simplified stochastic model of the puck dynamics for i) estimating the current state of the puck online, ii) predicting the trajectory of the puck online, and iii) solving a stochastic optimal control problem to plan the next best contact between the robot and the puck.
@@ III-A Mixture of linear-Gaussian Contact Dynamics
Suppose that $\bm{x}_{k}^{p}\in\mathbb{R}^{2}$ is the position of the puck w.r.t. the surface of the air hockey table at time step $k$. The robot interacts with the puck by making contact with its mallet, i.e. the circular part of the robot’s end-effector. The position of the mallet is denoted with $\bm{x}_{k}^{m}\in\mathbb{R}^{2}$ w.r.t. the surface of the air hockey table. We assume that the robot arm is controlled such that the mallet maintains contact with the table at all times. In order to efficiently perform rollouts of the puck dynamics, we impose a piecewise-linear structure on the model. Fig. 3 illustrates the three modes that we present in the following: 1) Floating, 2) Puck-Wall Collision, and 3) Puck-Mallet Collision. To account for modeling errors introduced through the piecewise-linear structure, we model each mode as a conditional Gaussian distribution, resulting in a mixture of linear-Gaussian contact dynamics. In the following, we present the individual modes and their respective parameters that are learned subsequently.
Figure 3: Illustration of three modes of the puck dynamics that are parameterized as linear-Gaussian modes. Mode 1) captures the dynamics of the puck $\bullet$ when floating on the surface of the table. Mode 2) captures collisions between puck and walls. Mode 3) models collisions between puck and the mallet $\bullet$ in a contact-aligned frame $\mathcal{C}$. The parameters for the nominal dynamics and the corresponding uncertainty are learned from data.
@@ III-B Learning Model Parameters from Data
Given the parameters of the joint probability distribution, the parameters of the linear-Gaussian models can be computed by conditioning the probability distribution on the input $\bm{\xi}$. Thus, the parameters are given by
@@ III-D Probability of Hitting the Goal
Fig. 4 illustrates stochastic rollouts for various initial conditions of a shot. Evaluating $\hat{\mathrm{G}}$ as defined in (11), we observe that those initial conditions have a significant effect even if the expected puck trajectory hits the center of the goal for all conditions. Fast shots (Fig. 4-b) accumulate less uncertainty compared to slow shots (Fig. 4-a) since the modeled process noise is constant over time. Direct shots accumulate less uncertainty during rollout than bank shots, as collisions with a wall add significant process noise (Fig. 4-c).
IV Implicit Contact Planning under Uncertainty
@@
The learned dynamics model enables the prediction of uncertain puck trajectories for contact planning. In particular, we search for contact states of the mallet that result in desired puck trajectories after contact. The proposed contact planning module is based on stochastic optimal control, optimizing the mallet’s state at contact. We combine this optimization with a model-based robot controller driving the robot to the desired contact state at the desired time (cf. Fig. 2).
@@ IV-B Shooting Angle as Reduced Action Space
IV-C Distilling an Optimal Contact Policy
The long-horizon predictions required to plan shooting actions make it difficult to operate at a rate that is sufficient for agile behavior. We address this challenge by training an implicit model to generate solutions to the stochastic optimal shooting problem (12) in realtime. Since the stochastic policy we aim to capture is multimodal, we opt for an implicit policy representation due to its ability to learn multimodal action distributions [8, 6]. In the case of air hockey, example modes correspond to the number of reflections against the bank of the table, ranging from straight shots without reflections to bank shots with multiple reflections. To learn the implicit policy, we use an energy-based model (EBM) [8]. Access to the learned energy [^1] In this paper, the term energy refers to the negative logarithm of the unnormalized probability density function that we want to model. landscape allows us to accommodate various sampling strategies without retraining, enabling a balance between optimality and computational efficiency.
We train the EBM by solving the computationally expensive shooting angle optimization offline and using the results as training data. Namely, we first generate a dataset of shooting angles for $N$ different puck states at time of contact $\{\bm{s}_{i}\}_{i=1}^{N}$. Due to the one-dimensional parametrization of the action space introduced in Sec. IV-B, we efficiently explore the space of shooting angles for each initial puck state by sampling $M$ candidate shooting angles $\{u_{i}^{j}\}_{j=1}^{M}$. Subsequently, we compute the stochastic rollout of the puck trajectory as presented in Sec. III-D and evaluate the objective and chance constraint from (12). The best-performing sample $\hat{u}_{i}$ is then used as a positive example for training the implicit behavioral cloning model, with the remaining $M-1$ samples as negative counter-examples. This results in a dataset of $M\times N$ state-action pairs $\{\bm{s}_{i},\hat{u}_{i},\{u_{i}^{j}\}_{j=1}^{M-1}\}_{i=1}^{N}$, which we use to train the EBM $E_{\theta}(\bm{s},u)$ using an InfoNCE-style [30] loss
$\mathcal{L}_{\text{InfoNCE}}$ $=\sum_{i=1}^{N}-\log\left(\tilde{p}_{\theta}\big{(}\hat{u}_{i}|\bm{s}_{i},\{u^{j}_{i}\}_{j=1}^{M-1}\big{)}\right).$ (14)
In the above, counter-examples $\{u^{j}_{i}\}_{j=1}^{M-1}$ are used to compute the likelihood $\tilde{p}_{\theta}$ as follows:
@@ IV-D Online Inference with Warm-Starting
To solve (12) given the estimated puck state $\hat{\bm{s}}_{0^{-}}$, we search for a state-action pair that minimizes the learned energy, i.e.
@@
For realtime optimization, we leverage direct access to the learned energy landscape of the EBM by executing sampling iterations concurrently with other components of the control loop. This allows us to refine contact plans as the robot executes trajectories, while considering only the shrinking horizon from the current timestep to the time of contact. We base the online retrieval of optimal shooting angles on derivative-free optimization procedures from [8]. As in the offline scenario, shooting actions are generated online by uniformly sampling $N$ candidate actions, inferring their energy values, and resampling with replacement to warm-start optimization in each subsequent timestep. To converge towards a solution with minimum implicit energy, reductions to the sampling variance are applied at each timestep, while keeping the optimal contact angle $\hat{u}$ as a reference for the mid-level trajectory planner. A full iteration of EBM inference is outlined in Alg. 1. We observe that the learned energy models and utilized optimization procedure efficiently retrieve multimodal contact plans to produce desired behaviors, as shown in Fig. 5.
@@ V-A2 EBM Architecture and Training
The energy-based shooting model consists of a multilayer perceptron with $2$ hidden, fully connected layers of $128$ neurons each. The model is trained on $N\!=\!3000$ initial puck states, with $M\!=\!100$ action samples. For each action sample, the corresponding state action pair is evaluated using the stochastic dynamics model learned from simulated data (Sec. III), and the optimal control cost (Sec. IV). Training required $500$ epochs to converge for satisfactory performance using the Adam [16] optimizer with a decaying learning rate.
@@ V-B Experimental Setup
Figure 6: The automated experimental setup consists of a robot placing the puck at a pre-defined grid of positions (left image). After releasing the puck, the KUKA robot executes the shooting policy (right image). We measure the speed of the puck at the goal line and whether the puck hit the goal as quality metrics.
The experiment is conducted with a KUKA iiwa14 LBR manipulator equipped with a mallet end-effector that is attached to a passive joint for seamless contact with the table surface. Experiments are carried out in a simulated MuJoCo environment and on a real-world setup (cf. Fig. 6) for evaluation of sim2real transfer. We evaluate three instances of the proposed approach by using different parameters for the chance-constrained optimization problem in (12). Ours #1: a conservative policy that prioritizes accuracy ($\lambda_{1}\!=\!1,\lambda_{2}\!=\!0,\beta\!=\!0.5$); Ours #2: a balanced policy that compromises between accuracy and puck speed ($\lambda_{1}\!=\!1,\lambda_{2}\!=\!0.2,\beta\!=\!0.5$); and Ours #3: an aggressive policy that prioritizes puck speed ($\lambda_{1}\!=\!0,\lambda_{2}\!=\!1,\beta\!=\!0.5$). Simulated results are also illustrated in Fig. 5 for initial puck velocities of zero. We compare the three instances of our contact planner with: CB, a baseline that utilizes conventional planning and control methods [18]; and ATACOM, a safe reinforcement learning approach for learning a robot policy [19]. Note that the ATACOM policy is trained in simulation and deployed in the physical experiment without additional tuning or retraining.
We perform 100 shots with each policy and report the accuracy score, puck speed at the goal line, and the number of bank reflections for successful shots. Each shot is initialized by placing the puck within a grid in front of the robot. Due to imperfect air flow on the air hockey table, the puck moves after release, requiring the robot to adapt for a good shot.
V-C Results
Recorded metrics are reported in Table III for the simulated environment and in Table II for the real-world environment. Compared to CB and ATACOM, we observe that our framework is capable of achieving higher scoring accuracy and significantly higher puck speeds in both environments. It can be seen that different instances of our policy obtain either a high score or high puck speeds according to the corresponding parameters of the stochastic optimal control problem. For example, when compared to Ours #1, it can be seen that Ours #3 compromises scoring accuracy for faster puck speeds and a high number of bank reflections, potentially making the shots more difficult to defend against. The higher number of bank reflections produced by Ours #3 indicates that the robot kinematics allow for higher shooting speeds when hitting laterally, at the risk of missing the goal due to uncertainty gained with every bank reflection. Note that the score of Ours #3 is lower than the score chance threshold $\beta=0.5$ for this instance. This indicates that the learned model either has an error in the nominal dynamics or expects too little uncertainty gain due to bank reflections. We further note a decrease in performance for all agents due to the sim2real gap, with ATACOM showing the highest sensitivity to transfer as it requires fine-tuning on the real environment. CB shows the least decrease in performance, as it is parameterized for the real system. However, note that the shooting trajectory optimization loop of CB is slower than required to run at 50 Hz, making it prone to errors due to the puck moving unpredictably during the shooting motion. Since we combine closed-loop model-based control with distilled contact planning, our agents display robustness to sim2real transfer. Additionally, we note higher puck speeds in the real setting for all agents as a result of differences in real and simulated contact dynamics. Example physical shots of all approaches can be found in the supplementary video.